# Feature Coding in the Era of Large Models: Dataset, Test Conditions, and Benchmark

### v1

@@ 4.1.3 Source data collection
Inspired by the video coding standard VVC [2], where test sequences are divided into 8 classes with 3–5 video sequences each, we organize our feature test dataset into five categories. Each category is supported by a curated subset of public datasets used as source data.
For image classification, we select 100 images from ImageNet [8], each representing a unique class accurately classified by DINOv2. For semantic segmentation, we collect 20 images from VOC2012, ensuring coverage of all 20 object classes. For depth estimation, we source 16 images from NYU-Depth-v2, adhering to the train-test split proposed in [21] and selecting one image for each scene category. In the common sense reasoning task, we utilize 100 samples from Arc-Challenge dataset, chosen for their longest input prompts and correct prediction by Llama3. For text-to-image synthesis, we first select 100 images from COCO2017 and then collect their longest captions.
@@ 4.2 Feature analysis
In this section, we analyze and compare the proposed features with those commonly used in existing research to clarify the necessity and importance of proposing a new dataset.
@@ 4.2.1 Distribution analysis
Frequency distributions and cumulative distribution functions (CDFs) of the proposed and commonly used existing features are visualized in Fig. 2. Features $C5$ and $C5$-ReLU are derived from Pytorch official ResNet50 pre-train on ImageNet, while $P$ features (including $p2$, $p3$, $p4$, and $p5$) are generated by the ResNeXT101-based MaskRCNN pre-trained on COCO2017.
The proposed features exhibit distinct distributions compared to existing features, underscoring the necessity and value of the dataset. For multi-layer split points, four key distinctions between $SP_{DM}$ and $P$ are observed: (1) $SP_{DM}$ distribution range expands progressively with deeper layers; (2) $SP_{DM}$ distribution is more asymmetric; (3) $SP_{DM}$ feature values are more concentrated, as shown in the CDFs; and (4) $SP_{DM}$ features contain more regions with sparse feature points. For single-layer split points, $SP_{GS}$ has a more concentrated distribution, while $SP_{H}$ shows more peaks. Unlike $C5$-ReLU, the proposed dataset omits similar features, as ReLU is infrequently used in large models.
The proposed features demonstrate the high diversity of the dataset, enhancing its representativeness and suitability of long-term research.
@@
TTI VAE Decoder CLIP Score
Table 2: Summary of task head and accuracy metric settings. Specific definitions of Heads can be found in the original papers [28, 11, 12].
@@ 4.2.2 Redundancy analysis
We conduct a redundancy analysis on feature blocks, using $7\times 7$ blocks for $C5$ and $C5$-ReLU and $16\times 16$ blocks for other features. We visualize the original feature blocks and their DCT blocks (absolute values) in Fig. 3. In the feature domain, clear spatial redundancy is observed in both $SP_{DM}$ and $SP_{H}$ features, with $SP_{H}$ exhibiting redundancy in both horizontal and vertical directions, while $SP_{DM}$ shows stronger vertical redundancy. By contrast, $SP_{GS}$ features vary rapidly in both directions, resulting in weak spatial redundancy. This is because Llama3 only processes textual data, which lacks inherent spatial correlation. In comparison, existing features exhibit smoother spatial variation and thus high spatial redundancy. We attribute this to the translation invariance of convolutional layers, which preserve the relative positions of pixels in the feature domain, resulting in consistent redundancy patterns. In contrast, transformer-based vision models partition images into patches before transforming them, shifting spatial redundancy primarily to the vertical direction, as seen in $SP_{DM}$.
In the DCT domain, the spatial redundancy is reflected by energy distribution: the proposed features display more distracted energy, whereas existing features concentrate energy in the top-left DC component. Notably, $SP_{H}$ has the highest energy concentration, as it looks similar to images.
The distinct redundancy characteristics indicate different requirements for feature coding methods, which should be explored in future studies.
Figure 4: The pipeline of the proposed baseline feature coding method. The quantization and de-quantization modules are only used in the VTM baseline.
@@
Metric BPFE Accuracy MSE BPFE mIoU MSE BPFE RMSE MSE BPFE Accuracy MSE BPFE CLIP Score MSE
Quantization 10 100 2.8974 10 79.93 1.8135 10 0.4941 0.5737 10 100 0.0042 10 30.07 0.00001
VTM Baseline 1.90 100 3.0117 1.78 79.68 1.9279 1.67 0.5341 0.6103 2.69 99 0.0114 1.30 30.04 0.0029
0.98 99 3.2324 0.90 78.91 2.1438 0.79 0.6925 0.6763 1.83 100 0.0267 0.65 29.95 0.0078
0.21 81 3.7424 0.24 73.53 2.6032 0.24 1.0684 0.8033 0.88 98 0.0826 0.26 29.50 0.0171
0.04 18 4.0751 0.04 55.05 3.0304 0.05 1.4053 0.9302 0.16 81 0.1900 0.10 27.43 0.0290
0.01 2 4.4588 0.01 32.64 3.3641 0.01 1.5542 1.0313 0.04 26 0.2483 0.04 24.42 0.0436
Hyperprior Baseline 1.99 92 3.5279 1.72 78.44 2.2636 1.51 0.5691 0.6588 5.50 95 0.0674 1.40 30.20 0.0054
1.11 91 3.7724 1.30 77.85 2.3423 1.01 0.6410 0.6830 2.84 89 0.0796 0.62 29.55 0.0122
0.89 86 3.9035 0.54 76.39 2.6461 0.43 0.9442 0.7818 1.46 83 0.1474 0.26 28.21 0.0216
0.37 29 4.3309 0.12 62.69 3.0935 0.08 1.1809 1.0830 1.30 67 0.1337 0.14 26.68 0.0294
0.23 11 4.8063 0.03 40.92 3.6286 0.01 2.5775 1.1876 1.19 32 0.1582 0.07 24.27 0.0404
Table 3: Rate-accuracy performance evaluations on the two baseline methods (MSE is calculated between $F_{o}$ and $\hat{F}_{r}$).
@@ 5.1 Bitrate computation
We introduce a new bitrate measurement, Bits Per Feature Point (BPFP), replacing the commonly used Bits Per Pixel (BPP). BPFP is calculated by dividing the total coding bits by the number of feature points.
Our rationale for shifting from BPP to BPFP is twofold. First, BPP is unsuitable for features extracted from non-visual data where pixels do not exist. Second, we believe bitrate should reflect the actual encoded data (the feature itself) rather than the source data. Using BPP to calculate bitrate on source data introduces ambiguity and limits fair comparisons between coding methods. For instance, applying the same BPP to different features extracted from the same image can be misleading. In addition, BPP may incentivize downscaling source images to reduce feature size, yielding lower bitrates that do not represent true coding efficiency. In contrast, BPFP measures bitrate on the feature itself, enabling fair comparisons across coding methods as long as the same feature is used.
@@ 5.2 Task accuracy evaluation
To isolate the impact of task heads on task accuracy, we fix their parameters during evaluation. All task head weights are loaded from pre-trained models. The task heads and accuracy metrics are presented in Table 2.
@@
6 The proposed baselines and benchmark
@@ 6.1 The baseline methods
The pipeline for the proposed feature coding methods is illustrated in Fig. 4. First, the original feature $F_{o}$ is truncated to a smaller range. The quantization operation is performed accordingly. Next, the quantized $F_{q}$ is packed into a 2D feature $F_{p}$. The encoder takes $F_{p}$ as input and outputs decoded feature $\hat{F}_{p}$, which is then unpacked to $\hat{F}_{q}$. $\hat{F}_{q}$ is converted back into a floating-point feature $\hat{F}_{r}$. We propose two baselines: a traditional coding method based on VTM and a learned coding method using the Hyperprior approach.
@@ 6.1.1 Pre-processing and post-processing
We propose two pre-processing operations: truncation and quantization. The truncation operation is designed to remove feature points that deviate significantly from the central region. The specific truncated regions for each task are listed in Table 4. Please note that different truncations are used for the two baselines for their distinct coding strategies. The quantization operation uniformly quantizes the truncated features to 10-bit integers. The post-processing only performs de-quantization, which is an inverse process of quantization.
Figure 5: Rate-accuracy-drop comparisons among different tasks. Left: VTM baseline. Right: Hyperprior baseline.
6.1.2 Packing and unpacking
Given the input requirement of VTM, we pack the features into a 2D YUV-400 format. For the Hyperprior encoder, we take the same packing to align with the VTM encoder. $F_{Seg}$, $F_{Dpt}$, and $F_{TTI}$ are packed into the shapes of $2740\times 1536$, $3222\times 6144$, and $512\times 512$, respectively. Please refer to supplementary materials for the packing details.
@@ 6.1.3 Encoding and decoding
For the VTM baseline, we employ the VVC reference software VTM-23.3 for feature coding. The Intra coding is applied using the main configuration file encoder_intra_vtm.cfg. The InputChromaFormat is set to 400, and ConformanceWindowMode is enabled. The InternalBitDepth, InputBitDepth, and OutputBitDepth are all set to 10. All other configurations are set to default. In our experiments, five QPs (22, 27, 32, 37, 42) are used.
For the Hyperprior baseline, we modify the input and output to accommodate the 2D features. In addition, we normalized all truncated features into a consistent region of [0,1] before training. Please find detailed training details in the supplementary.
6.2 Rate-accuracy analysis
The rate-accuracy evaluation results are presented in Table 3. We also provide the rate-accuracy curves in the supplementary materials. To further assess the effect of bitrate on accuracy, we introduce a new metric, accuracy drop, defined as the percentage decrease in accuracy. For depth estimation, the reciprocal of RMSE is employed. The bitrate-accuracy-drop curves, shown in Fig. 5, reveal that most tasks exhibit an inflection point where accuracy starts to decrease significantly. However, for depth estimation, this point is observed only in the Hyperprior baseline. In the Hyperprior baseline, inflection points are more distracted across tasks, which may caused by the fact that different models are trained for different tasks. Unlike other tasks, text-to-image synthesis shows smaller accuracy degradation even at low bitrates.
@@ 6.2 Rate-accuracy analysis
VTM Baseline 0.8933 0.8787 0.9904 0.7609 0.9329
Hyperprior Baseline 0.9190 0.9367 0.7741 0.5834 0.9793
Table 5: Linear correlation between MSE and task accuracy.

### v3

@@ 4.1.3 Source data collection
We organize our feature test dataset into five classes, each supported by a curated subset of publicly available datasets used as source data. For Cls, we select 500 images from ImageNet [8], each representing a unique class accurately classified by DINOv2. For Seg, we collect 100 images from VOC2012, ensuring coverage of all 20 object classes. For Dpt, we source 80 images from NYU-Depth-v2, adhering to the train-test split proposed in [22] and selecting 3-8 images for each scene category. For CSR, we utilize 500 samples from the Arc-Challenge dataset, chosen for their longest input prompts and correct prediction by Llama3. For TTI, we first select 500 images from COCO2017 and then collect their longest captions.
We noticed that learning-based methods may require a dedicated training set. We will provide the source code for generating features from any source data to serve as training data. Please find out our training data in the supplementary.
@@ 4.2 Feature analysis
In this section, we compare the proposed features with those commonly used in existing research to demonstrate the necessity and importance of introducing a new dataset.
@@ 4.2.1 Distribution analysis
We first compare frequency distributions and cumulative distribution functions (CDFs) of the proposed and commonly used existing features in Fig. 2. $C5$ and $C5$-ReLU are derived from the official ResNet50 pre-trained on ImageNet, while $P$-layer features are generated by the ResNeXT101-based MaskRCNN pre-trained on COCO2017.
For multi-layer split points, we observe four key distinctions between $SP_{DM}$ and $P$-layer features: (1) $SP_{DM}$ distribution range expands progressively with deeper layers; (2) $SP_{DM}$ distribution is more asymmetric; (3) $SP_{DM}$ feature values are more concentrated, as evidenced by CDFs; and (4) $SP_{DM}$ features contain more regions with sparse feature points. For single-layer split points, $SP_{GS}$ has a more concentrated distribution, while $SP_{H}$ shows more peaks. The proposed dataset omits $C5$-ReLU-like features, as ReLU is rarely used in large models.
The proposed features exhibit different distributions compared to the existing features, highlighting the necessity of introducing the new dataset. Additionally, the high diversity of the proposed features enhances the dataset’s representativeness and suitability for long-term research.
@@
TTI VAE Decoder FID
Table 2: Summary of task head and accuracy metric settings. Find definitions of Heads in the original papers [31, 11, 12].
@@ 4.2.2 Redundancy analysis
Since coding aims to remove redundancy in inputs, we conduct a redundancy comparison on feature blocks. We visualize the original feature blocks and their DCT blocks (absolute values) in Fig. 3. In the feature domain, clear spatial redundancy is observed in both $SP_{DM}$ and $SP_{H}$ features. $SP_{H}$ exhibits redundancy in both horizontal and vertical directions, while $SP_{DM}$ shows stronger vertical redundancy. In contrast, $SP_{GS}$ features vary rapidly in both directions, resulting in weak spatial redundancy. This is because Llama3 only processes textual data, which lacks inherent spatial correlation. Existing features, on the other hand, exhibit smoother spatial variation and higher spatial redundancy. We attribute this to the translation invariance of convolutional layers, which preserve the relative positions of pixels in the feature domain, resulting in consistent redundancy patterns. In contrast, transformer-based vision models partition images into patches before transforming them, shifting spatial redundancy primarily to the vertical direction, as seen in $SP_{DM}$.
In the DCT domain, spatial redundancy is reflected in the energy distribution. The proposed features display more dispersed energy, whereas existing features concentrate energy in the top-left DC component. $SP_{H}$ has the highest energy concentration as it closely resembles images.
These distinct redundancy characteristics further demonstrate the necessity of the proposed dataset and suggest that future studies should explore feature coding methods tailored to these unique properties.
Figure 4: The pipeline of the proposed baseline methods.
@@
Metric BPFP Accuracy MSE BPFP mIoU MSE BPFP RMSE MSE BPFP Accuracy MSE BPFP FID MSE
Original 32 100 0 32 83.39 0 32 0.3695 0 32 100 0 32 0 0
VTM Baseline 1.94 99.80 2.9499 1.76 80.42 1.8668 2.21 0.5021 0.6108 2.69 99.80 0.0109 1.41 6.14 0.0066
1.03 97.40 3.1693 0.88 79.31 2.0836 1.27 0.6511 0.6826 1.83 100 0.0260 0.74 19.94 0.0184
0.23 74.60 3.6993 0.23 73.47 2.5440 0.33 0.9530 0.9225 0.89 98.80 0.0810 0.29 58.42 0.0421
0.04 25.80 4.0651 0.04 56.53 2.9526 0.03 1.4850 1.0631 0.16 82.20 0.1868 0.11 109.23 0.0715
0.01 7.20 4.4400 0.01 37.42 3.2762 0.003 2.1174 1.1072 0.04 23.80 0.2455 0.05 171.63 0.1072
Hyperprior Baseline 2.01 93.20 3.4501 1.71 77.96 2.2010 1.51 0.4256 0.6600 6.34 91.40 0.0776 1.44 25.25 0.0138
1.13 88.80 3.6966 1.30 77.27 2.2795 1.01 0.4965 0.6844 3.60 87.80 0.0808 0.66 52.12 0.0300
0.91 83.60 3.8327 0.54 74.82 2.5845 0.43 0.6699 0.7837 1.68 82.40 0.1622 0.28 95.03 0.0541
0.37 29.00 4.2909 0.12 62.58 3.0151 0.08 1.0351 1.0844 1.50 57.80 0.1355 0.15 124.84 0.0734
0.23 14.60 4.7814 0.03 37.11 3.5449 0.01 2.7302 1.1866 1.35 34.80 0.1624 0.08 170.01 0.1006
Table 3: Rate-accuracy (R-A) performance evaluations on the two baseline methods (MSE is calculated between $F_{o}$ and $\hat{F}_{r}$).
@@ 5.1 Bitrate computation
We introduce a new bitrate measurement, Bits Per Feature Point (BPFP), replacing the commonly used Bits Per Pixel (BPP). BPFP is calculated by dividing the total coding bits by the number of feature points. Our rationale for shifting from BPP to BPFP is twofold. First, BPP is unsuitable for features extracted from non-visual data where pixels do not exist. Second, we believe bitrate should reflect the actual encoded data (the feature itself) rather than the source data. Using BPP to calculate bitrate on source data introduces ambiguity and limits fair comparisons between coding methods. For instance, applying the same BPP to different features extracted from the same image can be misleading. In addition, BPP may incentivize downscaling source images to reduce feature size, yielding lower bitrates that do not represent true coding efficiency. In contrast, BPFP measures bitrate on the feature itself, enabling fair comparisons across coding methods as long as the same feature is used.
@@ 5.2 Task accuracy evaluation
The decoded features undergo task heads to perform accuracy evaluation. To isolate the impact of task heads, we choose simple task heads and fix their parameters during evaluation. The task heads and accuracy metrics are presented in Table 2.
@@
6 Baselines and benchmark
As a starting point, we introduce two baselines and establish a benchmark. We adapt these methods to account for the characteristics of large model features. Our goal is to examine their applicability to large model feature coding and provide insights for interested researchers, rather than to develop a completely new method.
@@ 6.1 The baseline methods
The pipeline for the proposed baselines, illustrated in Fig. 4, includes pre-processing, core codec, and post-processing modules. First, the original feature $F_{o}$ is truncated to a smaller range and then quantized into integers. Next, the quantized $F_{q}$ is packed into a 2D feature $F_{p}$. The codec takes pre-processed $F_{p}$ as the input and outputs decoded feature $\hat{F}_{p}$. Two core codecs are used: a handcrafted codec VTM and a learning-based codec Hyperprior. In the post-processing module, $\hat{F}_{p}$ is unpacked to $\hat{F}_{q}$ and then converted back into a floating-point feature $\hat{F}_{r}$.
@@ 6.1.1 Pre-processing and post-processing
The truncation operation is proposed to remove feature points that deviate significantly from the central region. The specific truncated regions for each task are listed in Table 4. Different truncations are used for the two baselines for their distinct coding strategies. The quantization operation uniformly quantizes the truncated features to 10-bit integers. Given the input requirement of VTM, we pack the features into a 2D YUV-400 format. The same packing applies in the Hyperprior baseline. $F_{Seg}$, $F_{Dpt}$, and $F_{TTI}$ are packed into the shapes of $2740\times 1536$, $3222\times 6144$, and $512\times 512$, respectively. Please refer to supplementary materials for the packing details. The post-processing performs unpacking and de-quantization, which are inverse processes of packing and quantization.
Figure 5: Rate-accuracy-drop (R-AD) comparisons among different tasks. Left: VTM baseline. Right: Hyperprior baseline.
@@ 6.1.2 Encoding and decoding
VTM baseline: We employ VTM-23.3 as the codec. The Intra coding is applied using the main configuration file encoder_intra_vtm.cfg. The InputChromaFormat is set to 400, and ConformanceWindowMode is enabled. The InternalBitDepth, InputBitDepth, and OutputBitDepth are all set to 10. All other configurations are set to default. In our experiments, five QPs {22, 27, 32, 37, 42} are used.
Hyperprior baseline: We modify the channel of the input and output of the original Hyperprior as 1 to accommodate 2D features. The loss function is defined as:
$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
where $\lambda$ is a scaling factor used to adjust the bitrate. Please find more details in the supplementary.
6.2 Rate-accuracy (R-A) analysis
The R-A evaluation results are presented in Table 3. Both baselines can achieve significant bitrate savings, while the VTM baseline outperforms the Hyperprior baseline. To further assess the impact of bitrate on accuracy, we introduce a new metric, accuracy drop (AD), defined as the percentage decrease in accuracy. For Dpt, the reciprocal of RMSE is employed. The R-AD curves, shown in Fig. 5, reveal that most tasks exhibit an inflection point where accuracy starts to decrease significantly. However, for Dpt, this point is observed only in the Hyperprior baseline. In the Hyperprior baseline, inflection points are more distracted across tasks, which may caused by the fact that different models are trained for different tasks. TTI is excluded as images generated from the original features have an FID of 0.
6.3 Distortion-accuracy (D-A) analysis
We evaluate the linear correlation between feature MSE (D) and task accuracy (A) and present coefficients of determination in Table 5. Among the baselines, only TTI exhibits high linear correlations. This is reasonable since TTI involves generating images, and the baselines employ image codecs. In contrast, CSR fails to achieve a high linear correlation, as its features are derived from textual data. Overall, the VTM baseline outperforms the Hyperprior baseline, demonstrating its superior adaptability to large model features. Our analysis reveals that feature MSE struggles to characterize semantic distortion, particularly for textual data. This finding suggests that feature MSE is not an effective metric for measuring semantic distortion.
6.4 Complexity analysis
We run the VTM baseline and Hyperprior baseline on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively. The VTM baseline exhibits encoding times ranging from a few seconds to several hours, while its decoding time remains within a few seconds. In contrast, the Hyperprior baseline exhibits comparable encoding and decoding times, most of which are on the order of milliseconds. Detailed results are provided in the supplementary. We note that runtime depends heavily on both hardware capabilities and implementation details. A well-optimized codec and high-performance hardware can significantly accelerate feature coding.
@@ 6.4 Complexity analysis
VTM Baseline 0.9413 0.8955 0.8549 0.7490 0.9983
Hyperprior Baseline 0.9296 0.9280 0.7437 0.4713 0.9996
Table 5: Distortion-accuracy (D-A) analysis on the two baselines.
Task Seg CSR TTI
Models BPFP mIoU BPFP Acc. BPFP FID
Trained on Seg 1.71 77.96 2.08 83.80 0.97 119.83
1.30 77.27 1.37 66.60 0.62 208.50
0.54 74.82 0.78 0.20 0.31 333.75
0.12 62.58 0.25 0.00 0.12 363.32
0.03 37.11 0.09 0.00 0.05 354.22
Trained on TTI 2.58 78.30 2.28 97.60 1.44 25.25
1.96 76.34 1.46 77.80 0.66 52.12
1.12 73.37 0.65 32.00 0.28 95.03
0.53 56.90 0.30 0.00 0.15 124.84
0.26 42.94 0.11 0.00 0.08 170.01
Trained on CSR 5.29 78.20 6.34 91.40 4.77 58.19
3.27 77.94 3.60 87.80 2.59 73.19
1.80 76.77 1.68 82.40 1.02 196.56
1.73 75.82 1.50 57.80 1.02 191.69
1.54 75.80 1.35 34.80 0.94 193.08
Table 6: Generalizability evaluation on Seg, CSR, and TTI tasks.
6.5 Generalizability analysis
We assess the generalizability of trained models by compressing features extracted from one task using models trained on a different task and evaluating their R-A performance. We select three tasks and present the results in Table 6. Overall, for a given task (along column), three kinds of models behave differently in both bitrate and accuracy. For example, models trained on CSR struggle to achieve low bitrates when applied to Seg and TTI, while models trained on Seg fail to achieve low FID in TTI. More analyses can be found in the supplementary. In practical applications, an effective compression model should be capable of handling diverse inputs. Therefore, a compression method with high generalizability is highly desirable.
6.6 Discussion
Large model applications require a balance of high accuracy, low complexity, and strong generalizability. However, both baselines struggle to maintain high accuracy at low bitrates, underscoring the need for more efficient feature coding strategies tailored to the unique characteristics of large model features. In addition, semantic distortion measurement remains a critical challenge in feature coding. Advancing semantic distortion metrics could further optimize codec performance and enhance overall coding efficiency.
The two baseline approaches present distinct advantages and drawbacks. The handcrafted approach aligns seamlessly with existing codecs but suffers from excessive runtime, limiting its real-time applicability. In contrast, the learning-based approach benefits from efficient GPU acceleration but lacks generalizability, as separate models must be trained for different features. Therefore, developing lightweight and generic codecs remains a promising research direction for feature coding.
7 Conclusion
In this paper, we highlight the importance of feature coding in large model deployments and introduce a new research area: large model feature coding. As the first step, we provide a test dataset, unified test conditions, baseline methods, and a benchmark. We examine the applicability of the two baselines and identify promising directions for future research. Our aim is to encourage collaboration between image/video coding and large model communities to drive the development of feature coding.
Supplementary Material
Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.15, 10
Table 7: Training details of the Hyperprior baseline.
8 Details of baseline methods
8.1 Packing details
8.1.1 Semantic segmentation
In semantic segmentation, an image is pre-processed into multiple $518\times 518$ patches by crop and resize operations. These patches are fed into DINOv2, which produces features of shape $1370\times 1536$. In the proposed dataset, each image is pre-processed into 2 patches, resulting in features with a shape of $2\times 1370\times 1536$. We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
8.1.2 Depth estimation
In depth estimation, an image is first horizontally flipped. Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2. For each image, DINOv2 generates $4\times 1611\times 1536$ features at the split point $SP_{DM}$. Thus, the shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$. $F_{Dpt}$ is packed in two steps. First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width. Second, the two $1611\times 6144$ features (original first and then flipped images) are vertically stacked, forming a $3222\times 6144$ feature.
8.1.3 Text-to-image synthesis
8.2 Training details
8.2.1 Training dataset construction
We construct our training dataset from public datasets. The training features are extracted in the same way as the test dataset. The source code will be made publicly available, and we encourage researchers to extract features from their own source data.
As mentioned in the main text, raw features consume a significant amount of storage. During training, we encountered storage challenges. For example, 20000 original $F_{Dpt}$ features occupy 1474 GB of storage. As the demand for additional training data grows, storage requirements could increase substantially, often exceeding the resources available to researchers in academic institutions. To address this issue, we crop the features into smaller dimensions to facilitate feature coding training. The specific cropping configurations are detailed in Table 7.
It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training. For large models, cropping features to accelerate training is impractical, as downstream models typically require complete semantic information. For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation. Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
After feature extraction, the same truncation and quantization operations are applied to the training dataset. For the Hyperprior baseline, all features are normalized to the range [0, 1] before training.
8.2.2 Hyperparameter Setting
The hyperparameters, such as the number of training epochs, initial learning rate, and $\lambda$ values, are provided in Table 7. The number of training epochs is determined based on the size of the training dataset. In our experiments, all training processes converge to a stable loss. Five different $\lambda$ values are used to achieve various bitrates.
Metric QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec.
VTM Baseline 22 267.20 0.23 22 2301.30 1.11 22 14169.36 5.15 22 336.27 0.24 22 107.98 0.16
27 275.02 0.22 27 2039.34 0.94 27 10862.76 3.54 27 423.84 0.29 27 87.33 0.15
32 124.40 0.15 32 1399.71 0.52 32 7776.30 2.33 32 362.98 0.24 32 62.11 0.15
37 28.00 0.13 37 407.30 0.32 37 1590.27 1.14 37 162.05 0.17 37 32.38 0.24
42 8.55 0.14 42 70.66 0.27 42 348.68 1.17 42 66.28 0.18 42 15.70 0.15
Metric $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec.
Hyperprior Baseline 0.01 0.04 0.06 0.015 0.3 0.38 0.12 1.57 2.05 10 0.07 0.12 0.2 0.03 0.03
0.0035 0.04 0.05 0.007 0.31 0.4 0.05 1.59 2.05 0.07 0.07 0.09 0.05 0.03 0.03
0.003 0.04 0.05 0.003 0.31 0.39 0.02 1.55 1.88 0.015 0.06 0.08 0.02 0.03 0.03
0.0017 0.05 0.06 0.001 0.29 0.35 0.005 1.54 1.84 0.0142 0.07 0.08 0.01 0.03 0.03
0.001 0.05 0.06 0.0005 0.29 0.36 0.001 1.57 1.91 0.01405 0.06 0.08 0.005 0.03 0.03
Table 8: Encoding and decoding complexity evaluations. “Enc.” and “Dec.” denote encoding time and decoding time, respectively. The running time is measured in seconds.
$\lambda$-BPFP
Semantic Segmentation
Common Sense Reasoning
Text-to-Image Synthesis
Figure 6: $\lambda$-BPFP curves and R-A curves of generalizability evaluation on Seg, CSR, and TTI tasks.
9 Additional experimental results
$SP_{DM1},QP=42$
$SP_{DM1},\lambda=0.001$
$SP_{DM2},QP=42$
$SP_{DM2},\lambda=0.001$
$SP_{DM3},QP=42$
$SP_{DM3},\lambda=0.001$
$SP_{DM4},QP=42$
$SP_{DM4},\lambda=0.001$
$SP_{GS},QP=32$
$SP_{GS},\lambda=0.01405$
$SP_{H},QP=42$
$SP_{H},\lambda=0.005$
9.1 Complexity analysis
We present the encoding and decoding complexity analyses in Table 8. The VTM baseline and Hyperprior baseline are evaluated on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively. It is important to note that the VTM baseline is a handcrafted codec and cannot be executed on a GPU.
For the VTM baseline, encoding time increases as QP decreases and feature resolution increases. The block-based VTM codec exhaustively examines multiple partitions and modes during encoding, making it significantly more time-consuming than decoding. VTM is one of the reference implementations of the VVC standard. In practical deployments, a more efficient VVC codec, such as VVEnc, can be used to accelerate encoding. For the Hyperprior baseline, the decoding complexity is slightly higher than the encoding complexity. Across different bitrates, the encoding and decoding complexities remain comparable.
9.2 Generalizability analysis
Here, we visualize the $\lambda$-BPFP curves and the rate-accuracy curves for the cross evaluation to provide a more intuitive presentation of the results. The $\lambda$-BPFP curves reveal that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties. From the rate-accuracy curves, we observe that coding models trained for one task struggle to achieve comparable performance when applied to another task. The generalizability of feature coding methods across multiple tasks has received limited attention in prior research. By presenting these results, we aim to highlight this important research direction and encourage further exploration.
9.3 Ablation on truncation and quantization
In this subsection, we examine the impact of truncation and quantization on feature coding. As shown in Table 10, truncation and quantization individually result in little task accuracy loss. When both are applied together, task accuracies experience slight decreases. However, applying the VTM baseline to quantized features leads to a significant reduction in accuracy.
Without truncation, feature values are often limited to a small subset within the full range of [0, 1023]. In this scenario, during feature coding, the codec’s internal quantization maps values with a small range to a single value, introducing significant distortion. By applying truncation, this range is expanded, which helps mitigate distortion caused by VTM’s internal quantization. To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37). As presented in Table 11, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates. Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
Proposed Features Existing Features
Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
Captions from COCO2017 $SP_{H}$ 3.67 -2.11 0.5664 21130.12 138.53 / / / / / /